Every human has blood type A, B, O or AB. 1/10 have type B and 23/50 have type O.

What is the probability that someone has blood type A or AB?

Answer by anand429(138)

Let total humans = 50x
No. of human having type B = (1/10)*50x = 5x
No. of human having type O = (23/50)*50x = 23x
So,
No. of human having type A or AB = 50x -5x -23x
= 22x
So probability that someone has blood type A or AB = 22x/50x
= 11/25